OpenStudy (anonymous):

. In solving the equation (x + 2)(x – 2) = 32, Eric stated that the solution would be x + 2 = 32 => x = 30 or (x – 2) = 32 => x = 34 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

Multiply out (x+2)(x-2)=32 x^2-2x +2x -4 =32 Then simplify x^2=36 x=6 You have to be careful of the signs